OpenStudy (anonymous):

perform the operation on complex numbers @Isaiah.Feynman @ilfy214 @seesawn

OpenStudy (anonymous):

\[(2+i)/(3+2i)\]

OpenStudy (anonymous):

Operation?

OpenStudy (anonymous):

like simplify i think

OpenStudy (anonymous):

Okay well try to get rid of your denominator multiplying your num and denim by (3+2i) \[\frac{ (2+i)}{(3+2i)}\] \[\frac{ (2+i)}{(3+2i)} \times \frac{ (3+2i)}{(3+2i)} \]

OpenStudy (anonymous):

what/

OpenStudy (anonymous):

how

OpenStudy (anonymous):

multiply? You know how

OpenStudy (anonymous):

i mean like am i cross multiplying

OpenStudy (anonymous):

Hold on. I meant THIS: \[\frac{ (2+i)}{(3+2i)} \times \frac{ (3-2i)}{(3-2i)}\]

OpenStudy (anonymous):

And you do not cancel anything out here

OpenStudy (anonymous):

i got 8/13-1/13

OpenStudy (anonymous):

@ilfy214

OpenStudy (anonymous):

almost... not 1! \[\frac{ 8}{13}-\frac{ i}{13}\]

OpenStudy (anonymous):

ooo ok